OpenStudy (anonymous):

write the ratio in lowest terms in order to decide whether the proportion is true or false. 10/15=5/8 what is the proportion in lowest terms?(type a fraction)

OpenStudy (anonymous):

\[\frac{10}{15}=\frac{2}{3}\]

OpenStudy (anonymous):

wrong=/

OpenStudy (espex):

10/15 = 5/8, write the them both in their lowest terms. Divide numerator and denominator by 5, you get 2/3. Now look at 5/8, 5 cannot be reduced therefore it is in its lowest terms, SO, 2/3 = 5/8, the answer is false.